# Robust iterative methods for linear systems with saddle point structure

### v2

@@ 1 Introduction
Typical applications where linear systems of the form (1) arise are in the time integration of dissipative Hamiltonian systems, see [17, 24], in computational fluid dynamics [7, 11], and in constrained optimization problems [12].
In most applications, one has $C=B$ but the case $B\neq C$ also arises in practice, see the examples presented in Section 4. Very often the two components of the solution vector represent different quantities, such as e.g. a state vector $x$ and a Lagrange multiplier $y$, that penalizes the constraints.
@@
    the ’generalized case’, where $B=C$ but $A\not=A^{T}$ and $A+A^{T}\geq 0$, and
@@ 2 Theoretical Analysis
Consider a symmetric, generalized or general linear system of the form (1), with $A+A^{T}\geq 0$, and $\mathop{\rm rank}\nolimits B=\mathop{\rm rank}\nolimits C^{T}=k\leq m$. Then there exist real orthogonal matrices $Q,V$ such that
@@
Moreover, in the symmetric and generalized case $B_{21},B_{31},B_{41}=0$.
@@
In the symmetric and generalized case we also have that $x_{1}=-C_{11}^{-T}g_{1}$, $x_{3}=A_{33}^{-1}(A_{31}C_{11}^{-T}g_{1}+f_{3})$, $x_{2}=A_{22}^{-1}(-A_{21}x_{1}+f_{2})$, and $x_{4}$ is arbitrary. So if $\mathcal{W}$ is invertible, then the forth row and column must be void as well.
@@
ii) For the generalized case, i.e., if $A\neq A^{T}$, but $A+A^{T}\geq 0$, $B=C$, and $\mathcal{W}$ is invertible, then there exists a real orthogonal matrix $Q$ such that $Q^{T}Z^{T}AZQ=\mbox{\rm diag}(A_{22},A_{33})$, with $A_{33}$ skew-symmetric and invertible. If the matrix $A$ can be transformed to the form (6) with the last two rows and columns void, then the block $A_{33}$ is void and for $A_{22}$ there exists a block $LDU$ factorization as in (7), where the last block is void and the symmetric part is positive definite.
@@
The construction of the projection matrices $Z,U$ with orthonormal columns will typically lead to full matrices. So for the practical implementation we will use sparse approximate versions of the associated decompositions. In this way we will not obtain exact bases for the nullspaces, but sparse approximations $\tilde{Z},\tilde{U}$. For the symmetric and generalized case, the following result shows that if the error $\|Z-\tilde{Z}\|$ can be bounded by a small quantity, then the approximate nullspace matrix $\tilde{Z}^{T}A\tilde{Z}$ will still have a positive definite symmetric part.
@@
Under the assumptions of Corollary 1, let in the symmetric and generalized case $\tilde{Z}$ be a (sparse) approximation of $Z$ (with columns that are not exactly orthonormal) such that $\|\tilde{Z}-Z\|_{2}\leq\epsilon\|Z\|_{2}=\epsilon$, with a sufficiently small $\epsilon$. Then $\tilde{Z}^{T}A\tilde{Z}$ still has a positive definite symmetric part.
@@ 3 A new multi-layer iterative scheme
While for the solution of saddle point problems in the symmetric and generalized case, nullspace methods have been studied extensively [28, 32], the general case has not received much attention, due to the associated computational and memory cost, which involves two different nullspace bases. Typically, furthermore, for the usual nullspace approaches it is required that the nullspace dimension is small. In contrast to this, by allowing sparse approximations, we show how these issues can be alleviated in our new technique.
@@
A pseudocode for the sparse approximate right oblique conjugation with pivoting to obtain a sparse approximate nullspace basis is presented in Algorithm 2. Hereafter, we refer to this procedure as SAROC. It produces a sparse basis, since we assume that computed coefficients that are smaller than a given threshold and entries of the matrix for which we compute the nullspace that are lower than a given tolerance in absolute value are set to zero, as shown in lines 8 and 9, respectively. As a consequence, we obtain approximate projected nullspace matrices, $\widetilde{\mathcal{N}}=\widetilde{Z}^{T}A\widetilde{Z}$ and $\widetilde{\mathcal{N}}=\widetilde{Z}^{T}A\widetilde{U}$, respectively, for the symmetric/generalized and general case.
@@ 3.1 The symmetric case
3.2 Generalized case
In the generalized case, we assume $B=C$ and that $A\neq A^{T}$. We compute a sparse approximate nullspace basis, $\widetilde{Z}$, of $B^{T}$. We again assume that $\widetilde{\mathcal{N}}^{s}=\widetilde{Z}^{T}(\frac{A+A^{T}}{2})\widetilde{Z}>0$, see Section 2. The pseudocode is given in Algorithm 6. We perform matrix vector multiplications with $\widetilde{\mathcal{N}}^{s}$ and $\widetilde{\mathcal{N}}^{j}$ implicitly, while forming the symmetric and skew-symmetric parts of $A$ explicitly.
@@ 3.2 Generalized case
Algorithm 6 New Scheme (generalized case)
@@ 3.3 General case
In the general case, we allow $B\neq C$ and $A\neq A^{T}$ and we compute sparse approximate nullspace bases, $\widetilde{Z}$ and $\widetilde{U}$ of $B^{T}$ and $C^{T}$, respectively. Here we again assume that the symmetric part $\widetilde{\mathcal{N}}^{s}=(\widetilde{Z}^{T}A\widetilde{U}+\widetilde{U}^{T}A^{T}\widetilde{Z})/2$, of $\widetilde{Z}^{T}A\widetilde{U}$, is positive definite. The pseudocode is given in Algorithm 7. Similar to the generalized case, in the general case, we also perform the vector multiplications with $\widetilde{\mathcal{N}}^{s}$ and $\widetilde{\mathcal{N}}^{j}$ implicitly, but this requires matrix vector multiplications involving both $A$ and $A^{T}$.
Unlike in the symmetric and generalized case, rather than $M$-orthogonalization, one needs to biorthogonalize $\widetilde{Z}$ and $\widetilde{U}$ with respect to $\widetilde{\mathcal{N}}^{s}$. One way to achieve this is by computing the FSAI of $\widetilde{\mathcal{N}}^{s}$. In other words, $\widetilde{Z}\widetilde{W}$ and $\widetilde{U}\widetilde{W}$ are already bi-orthogonalized with respect to $\widetilde{\mathcal{N}}^{s}$. We do not perform any further bi-orthgonalization step in the general case.
@@ 4 Test problems
For the generalized case, eight matrices are used (given in Table 2). The first three, namely rajat04, rajat14 and fpga_trans_01, are from the SuiteSparse Matrix Collection. The remaining five test matrices are generated from the IFISS software [11] for the steady-state solution of a linearized 2D Navier-Stokes equations and are obtained from for the lid-driven cavity problem with a spatial discretization based on $Q_{2}/Q_{1}$ elements and a uniform grid, varying the Reynolds number between $100$ and $900$. Since our method is a black-box algebraic approach, we ignore the fact that these problems arise from specific applications and treat them just as given test matrices. The $B$ matrix in this case is numerically rank-deficient. Therefore, in the new scheme, we compute a nullspace basis of size mn+1m-n+1. In Table 2, we provide the smallest and largest eigenvalues of $A+A^{T}$. For three of the test problems in this set, $A+A^{T}$ is positive definite but close to singular, while for the remaining five problems it is robustly positive definite.
@@
Table 2: Generalized case: matrix properties
@@ 5 Experimental framework
The residual norms are obtained as a by-product of fGMRES and GMRES. Here, right preconditioned fGMRES produces the true residual norm, and GMRES produces the preconditioned residual norm due to ILUTP-based left and right preconditioning. For these reasons, we also calculate and report the final true relative residual norms. For the new schemes, for the symmetric case, the inner LSQR and CG iterations are stopped when the $L_{2}$-norm of the relative residual is less than or equal to $\epsilon_{in}$ and $\epsilon_{innermost}$ is not used. For the generalized and general case, inner LSQR and fGMRES iterations are stopped when the $L_{2}$-norm of the relative residual is less than or equal to $\epsilon_{in}$ and the innermost MRS iterations are stopped when the $L_{2}$-norm of the relative residual is less than or equal to $\epsilon_{innermost}$. Since our new schemes are based on multiple nested loops and external functions, a fair comparison in terms of the running times against MATLAB’s built-in functions is currently not possible, and we only provide comparisons in terms of the number of iterations, number of nonzeros, and the robustness of the new and baseline schemes. The number of iterations and the number of nonzeros provide approximate measures for computational time and memory requirements.
@@ 8 $\ddagger$ 1 1
6.2 Generalized case
As before, let us first consider the new scheme without $M$-orthogonalization. For the problems given in Table 2, in Table 11 we present the number of outer GMRES(10) and fGMRES(10) iterations for the baseline method and the new scheme, respectively. Here, ILUTP frequently fails, as it encounters a zero pivot, sometimes even when a stringent or a relaxed drop tolerance is used but converges for some arbitrary values of drop tolerances for the lid-driven cavity problem. In a number of cases, ILUTP fails because the final true relative residual norm is larger than the stopping tolerance. The new method converges for all problems without any failure, showing the robustness of the new scheme for the generalized problem set.
@@ 6.2 Generalized case
Table 11: Number of outer fGMRES/GMRES iterations for the generalized case
@@
Table 12: Number of nonzeros of the preconditioners for the generalized case
In Table 13, we present the average number of inner fGMRES(10), LSQR and MRS iterations for the new scheme. While all are reasonably small for the circuit simulation systems, the number of inner fGMRES(10) and MRS iterations grow as the Reynold’s number increases for the lid-driven cavity problem. However, this growth is suppressed if $\tau_{large}$ is used. For generalized test problems, the number of LSQR iterations remained reasonably small for all problems, even without using any preconditioner. Figure 2 shows the final true relative residual norms of the new scheme and the baseline for the generalized case. The horizontal dashed line visualizes the stopping tolerance where the norm of the relative residual is equal to $10^{-5}$.
Figure 2: Final true relative residual norms of the new scheme and the baseline for the generalized case
@@
Table 13: Average number of inner fGMRES, LSQR and MRS iterations for the new scheme (generalized case)
Next, we consider the additional $M$-orthogonalization of the nullspace basis via the modified Gram-Schmidt procedure. The number of outer fGMRES(10) iterations and the total number of nonzeros for the new scheme are given in Tables 14 and 15, respectively. As before, the total number of nonzeros refers to the sum of the number of nonzeros of the nullspace basis, the $M$-orthogonalized nullspace basis and the nonzeros of the SPAI preconditioner (i.e. nnz(Z~)+nnz(Z¯)+nnz(W~)nnz($\widetilde$$\bar{Z})+nnz(\tilde${W})). The average number of inner fGMRES(10), LSQR, and MRS are given in Table 16. For test problems in the generalized case, we do not observe any significant differences when $M$-orthogonalization is performed. There is some increase in the number of nonzeros in most cases.
@@
Table 14: Number of outer fGMRES iterations for the new scheme (generalized case) with $M$-orthogonalization
@@
Table 15: Number of nonzeros of the new scheme for the new scheme (generalized case) with $M$-orthogonalization
@@
Table 16: Average number of inner fGMRES, LSQR and MRS iterations for the new scheme (generalized case) with $M$-orthogonalization
@@ 8 $114\ddagger$ 90 134 $\dagger$ $\dagger$ $\dagger$ $\dagger$
Figure 3: Final true relative residual norms of the new scheme and the baseline for the generalized case
@@ 7 Conclusions
We have introduced a new class of multilayer iterative schemes for solving sparse linear systems in saddle point structure. We have presented a theoretical analysis for the (approximate) nullspace method. We have demonstrated the effectiveness of the new schemes by solving linear systems that arise in a variety of applications and compare them to a classical preconditioned iterative scheme. The new schemes may benefit from further preconditioning in the Krylov subspace methods for solving over/underdetermined linear least squares problems. This is currently under investigation. Although rarely encountered in practice, we believe that a further generalization of the new schemes for linear systems where the $(2,2)$-block is non-square, as well as generalizing the new schemes for over/underdetermined linear least squares problems with saddle point structures is possible. Furthermore, a parallel implementation of the new schemes for large-scale problems is also planned as a future work.

### v3

# Multilayer approximate nullspace methods for saddle point systems

@@ Abstract
Abstract
We propose a new class of multi-layer iterative schemes for solving sparse linear systems in saddle point structure. The new scheme consists of an iterative preconditioner that is based on the (approximate) nullspace method, combined with an iterative least squares approach and an iterative projection method. We present a theoretical analysis and demonstrate the effectiveness and robustness of the new scheme on sparse matrices from various applications.
@@ 1 Introduction
Typical applications where linear systems of the form (1) arise are in the time integration of dissipative Hamiltonian systems, see [18, 25], in computational fluid dynamics [7, 11], and in constrained optimization problems [13]. In most applications, one has $C=B$ but the case $B\neq C$ also arises in practice, see the examples presented in Section 4. Very often the two components of the solution vector represent different quantities, such as e.g. a state vector $x$ and a Lagrange multiplier $y$, that penalizes the constraints.
@@
    the ’structurally symmetric case’, where $B=C$ but $A\not=A^{T}$ and $A+A^{T}\geq 0$, and
@@ 2 Theoretical Analysis
Consider a symmetric, structurally symmetric or general linear system of the form (1), with $A+A^{T}\geq 0$, and $\mathop{\rm rank}\nolimits B=\mathop{\rm rank}\nolimits C^{T}=k\leq m$. Then there exist real orthogonal matrices $Q,V$ such that
@@
Moreover, in the symmetric and structurally symmetric case $B_{21},B_{31},B_{41}=0$.
@@
In the symmetric and structurally symmetric case we also have that $x_{1}=-C_{11}^{-T}g_{1}$, $x_{3}=A_{33}^{-1}(A_{31}C_{11}^{-T}g_{1}+f_{3})$, $x_{2}=A_{22}^{-1}(-A_{21}x_{1}+f_{2})$, and $x_{4}$ is arbitrary. So if $\mathcal{W}$ is invertible, then the forth row and column must be void as well.
@@
ii) For the structurally symmetric case, i.e., if $A\neq A^{T}$, but $A+A^{T}\geq 0$, $B=C$, and $\mathcal{W}$ is invertible, then there exists a real orthogonal matrix $Q$ such that $Q^{T}Z^{T}AZQ=\mbox{\rm diag}(A_{22},A_{33})$, with $A_{33}$ skew-symmetric and invertible. If the matrix $A$ can be transformed to the form (6) with the last two rows and columns void, then the block $A_{33}$ is void and for $A_{22}$ there exists a block $LDU$ factorization as in (7), where the last block is void and the symmetric part is positive definite.
@@
The construction of the projection matrices $Z,U$ with orthonormal columns will typically lead to full matrices. So for the practical implementation we will use sparse approximate versions of the associated decompositions. In this way we will not obtain exact bases for the nullspaces, but sparse approximations $\tilde{Z},\tilde{U}$. For the symmetric and structurally symmetric case, the following result shows that if the error $\|Z-\tilde{Z}\|$ can be bounded by a small quantity, then the approximate nullspace matrix $\tilde{Z}^{T}A\tilde{Z}$ will still have a positive definite symmetric part.
@@
Under the assumptions of Corollary 1, let in the symmetric and structurally symmetric case $\tilde{Z}$ be a (sparse) approximation of $Z$ (with columns that are not exactly orthonormal) such that $\|\tilde{Z}-Z\|_{2}\leq\epsilon\|Z\|_{2}=\epsilon$, with a sufficiently small $\epsilon$. Then $\tilde{Z}^{T}A\tilde{Z}$ still has a positive definite symmetric part.
@@ 3 A new multi-layer iterative scheme
While for the solution of saddle point problems in the symmetric and structurally symmetric case, nullspace methods have been studied extensively [29, 33], the general case has not received much attention, due to the associated computational and memory cost, which involves two different nullspace bases. Typically, furthermore, for the usual nullspace approaches it is required that the nullspace dimension is small. In contrast to this, by allowing sparse approximations, we show how these issues can be alleviated in our new technique.
@@
A pseudocode for the sparse approximate right oblique conjugation with pivoting to obtain a sparse approximate nullspace basis is presented in Algorithm 2. Hereafter, we refer to this procedure as SAROC. It produces a sparse basis, since we assume that computed coefficients that are smaller than a given threshold and entries of the matrix for which we compute the nullspace that are lower than a given tolerance in absolute value are set to zero, as shown in lines 8 and 9, respectively. As a consequence, we obtain approximate projected nullspace matrices, $\widetilde{\mathcal{N}}=\widetilde{Z}^{T}A\widetilde{Z}$ and $\widetilde{\mathcal{N}}=\widetilde{Z}^{T}A\widetilde{U}$, respectively, for the symmetric/structurally symmetric and general case. Assuming dense operations, the worst-case cost of SAROC is $\mathcal{O}(n^{2})$ arithmetic operations per iteration, giving a total cost of $\mathcal{O}(kn^{2})$. However, since the matrix $V$ is initialized as the identity matrix and the algorithm is designed to maintain sparsity via the drop tolerance ($\tau$) and the threshold ($\rho$)-controlled vector updates, the cost is reduced significantly. Let $s_{v}$ and $s_{b}$ denote the average number of nonzeros per column of $V$ and $B$, respectively. Inner products (line 3) cost $\mathcal{O}(sn)$ per iteration and vector updates (line 9) cost $\mathcal{O}(s_{v})$ per update, where $s=\min(s_{v},s_{b})$. Since the threshold $\rho$ limits the updates (worst-case nin-i per iteration), the total cost would be approximately $\mathcal{O}(ksn)$ where $s\ll n$. Furthermore, since the algorithm performs column pivoting (lines 4-6), the matrix $V$ must be stored in column-major order to avoid cache misses.
@@ 3.1 Symmetric case
3.2 Structurally symmetric case
In the structurally symmetric case, we assume $B=C$ and that $A\neq A^{T}$. We compute a sparse approximate nullspace basis, $\widetilde{Z}$, of $B^{T}$. We again assume that $\widetilde{\mathcal{N}}^{s}=\widetilde{Z}^{T}(\frac{A+A^{T}}{2})\widetilde{Z}>0$, see Section 2. The pseudocode is given in Algorithm 6. We perform matrix vector multiplications with $\widetilde{\mathcal{N}}^{s}$ and $\widetilde{\mathcal{N}}^{j}$ implicitly, while forming the symmetric and skew-symmetric parts of $A$ explicitly.
@@ 3.2 Structurally symmetric case
Algorithm 6 New Scheme (structurally symmetric case)
@@ 3.3 General case
In the general case, we allow $B\neq C$ and $A\neq A^{T}$ and we compute sparse approximate nullspace bases, $\widetilde{Z}$ and $\widetilde{U}$ of $B^{T}$ and $C^{T}$, respectively. Here we again assume that the symmetric part $\widetilde{\mathcal{N}}^{s}=(\widetilde{Z}^{T}A\widetilde{U}+\widetilde{U}^{T}A^{T}\widetilde{Z})/2$, of $\widetilde{Z}^{T}A\widetilde{U}$, is positive definite. The pseudocode is given in Algorithm 7. Similar to the structurally symmetric case, in the general case, we also perform the vector multiplications with $\widetilde{\mathcal{N}}^{s}$ and $\widetilde{\mathcal{N}}^{j}$ implicitly, but this requires matrix vector multiplications involving both $A$ and $A^{T}$.
Unlike in the symmetric and structurally symmetric case, rather than $M$-orthogonalization, one needs to biorthogonalize $\widetilde{Z}$ and $\widetilde{U}$ with respect to $\widetilde{\mathcal{N}}^{s}$. One way to achieve this is by computing the FSAI of $\widetilde{\mathcal{N}}^{s}$. In other words, $\widetilde{Z}\widetilde{W}$ and $\widetilde{U}\widetilde{W}$ are already bi-orthogonalized with respect to $\widetilde{\mathcal{N}}^{s}$. We do not perform any further bi-orthgonalization step in the general case.
@@ 4 Test problems
For the structurally symmetric case, eight matrices are used (given in Table 2). The first three, namely rajat04, rajat14 and fpga_trans_01, are from the SuiteSparse Matrix Collection. The remaining five test matrices are generated from the IFISS software [11] for the steady-state solution of a linearized 2D Navier-Stokes equations and are obtained from for the lid-driven cavity problem with a spatial discretization based on $Q_{2}/Q_{1}$ elements and a uniform grid, varying the Reynolds number between $100$ and $900$. Since our method is a black-box algebraic approach, we ignore the fact that these problems arise from specific applications and treat them just as given test matrices. It is clear that problem specific solvers, see [12], that exploit the underlying discrete differential properties of the operators will, in general, show better performance than black-box methods. The $B$ matrix in this case is numerically rank-deficient. Therefore, in the new scheme, we compute a nullspace basis of size mn+1m-n+1. In Table 2, we provide the smallest and largest eigenvalues of $A+A^{T}$. For three of the test problems in this set, $A+A^{T}$ is positive definite but close to singular, while for the remaining five problems it is robustly positive definite.
@@
Table 2: Structurally symmetric case: matrix properties
@@ 5 Experimental framework
The residual norms are obtained as a by-product of fGMRES and GMRES. Here, right preconditioned fGMRES produces the true residual norm, and GMRES produces the preconditioned residual norm due to ILUTP-based left and right preconditioning. For these reasons, we also calculate and report the final true relative residual norms. For the new schemes, for the symmetric case, the inner LSQR and CG iterations are stopped when the $L_{2}$-norm of the relative residual is less than or equal to $\epsilon_{in}$ and $\epsilon_{innermost}$ is not used. For the structurally symmetric and general case, inner LSQR and fGMRES iterations are stopped when the $L_{2}$-norm of the relative residual is less than or equal to $\epsilon_{in}$ and the innermost MRS iterations are stopped when the $L_{2}$-norm of the relative residual is less than or equal to $\epsilon_{innermost}$. Since our new schemes are based on multiple nested loops and external functions, a fair comparison in terms of the running times against MATLAB’s built-in functions is currently not possible, and we only provide comparisons in terms of the number of iterations, number of nonzeros, and the robustness of the new and baseline schemes. The number of iterations and the number of nonzeros provide approximate measures for computational time and memory requirements.
@@ 8 $\ddagger$ 1 1
6.2 Structurally symmetric case
As before, let us first consider the new scheme without $M$-orthogonalization. For the problems given in Table 2, in Table 11 we present the number of outer GMRES(10) and fGMRES(10) iterations for the baseline method and the new scheme, respectively. Here, ILUTP frequently fails, as it encounters a zero pivot, sometimes even when a stringent or a relaxed drop tolerance is used but converges for some arbitrary values of drop tolerances for the lid-driven cavity problem. In a number of cases, ILUTP fails because the final true relative residual norm is larger than the stopping tolerance. The new method converges for all problems without any failure, showing the robustness of the new scheme for the structurally symmetric problem set.
@@ 6.2 Structurally symmetric case
Table 11: Number of outer fGMRES/GMRES iterations for the structurally symmetric case
@@
Table 12: Number of nonzeros of the preconditioners for the structurally symmetric case
In Table 13, we present the average number of inner fGMRES(10), LSQR and MRS iterations for the new scheme. While all are reasonably small for the circuit simulation systems, the number of inner fGMRES(10) and MRS iterations grow as the Reynold’s number increases for the lid-driven cavity problem. However, this growth is suppressed if $\tau_{large}$ is used. For structurally symmetric test problems, the number of LSQR iterations remained reasonably small for all problems, even without using any preconditioner. Figure 2 shows the final true relative residual norms of the new scheme and the baseline for the structurally symmetric case. The horizontal dashed line visualizes the stopping tolerance where the norm of the relative residual is equal to $10^{-5}$.
Figure 2: Final true relative residual norms of the new scheme and the baseline for the structurally symmetric case
@@
Table 13: Average number of inner fGMRES, LSQR and MRS iterations for the new scheme (structurally symmetric case)
Next, we consider the additional $M$-orthogonalization of the nullspace basis via the modified Gram-Schmidt procedure. The number of outer fGMRES(10) iterations and the total number of nonzeros for the new scheme are given in Tables 14 and 15, respectively. As before, the total number of nonzeros refers to the sum of the number of nonzeros of the nullspace basis, the $M$-orthogonalized nullspace basis and the nonzeros of the SPAI preconditioner (i.e. nnz(Z~)+nnz(Z¯)+nnz(W~)nnz($\widetilde$$\bar{Z})+nnz(\tilde${W})). The average number of inner fGMRES(10), LSQR, and MRS are given in Table 16. For test problems in the structurally symmetric case, we do not observe any significant differences when $M$-orthogonalization is performed. There is some increase in the number of nonzeros in most cases.
@@
Table 14: Number of outer fGMRES iterations for the new scheme (structurally symmetric case) with $M$-orthogonalization
@@
Table 15: Number of nonzeros of the new scheme for the new scheme (structurally symmetric case) with $M$-orthogonalization
@@
Table 16: Average number of inner fGMRES, LSQR and MRS iterations for the new scheme (structurally symmetric case) with $M$-orthogonalization
@@ 8 $114\ddagger$ 90 134 $\dagger$ $\dagger$ $\dagger$ $\dagger$
Figure 3: Final true relative residual norms of the new scheme and the baseline for the structurally symmetric case
@@ 7 Conclusions
We have introduced a new class of multilayer iterative schemes for solving sparse linear systems in saddle point structure. We have presented a theoretical analysis for the (approximate) nullspace method. We have demonstrated the effectiveness of the new schemes by solving linear systems that arise in a variety of applications and compare them to a classical preconditioned iterative scheme. The new schemes may benefit from further preconditioning in the Krylov subspace methods for solving over/underdetermined linear least squares problems. This is currently under investigation. Furthermore, a parallel implementation of the new schemes for large-scale problems is also planned as a future work.